Sequential Method

Eilers Company has two producing departments and two support departments. The following budgeted data pertain to these four departments:


Support Departments
Producing Departments
General
Factory
Receiving Assembly Finishing
Direct overhead $400,000 $160,000 $43,000 $74,000
Square footage 2,700 5,400 5,400
Number of receiving orders 300 1,680 1,020
Direct labor hours 25,000 40,000

The support departments are ranked in order of highest cost to lowest cost.

Required:

1. Allocate the costs of the support departments using the sequential method. (Use the rounded values for subsequent calculations. Round allocation ratios to four significant digits. Round allocated costs to the nearest dollar. If an amount is zero, enter "0".)

Allocation ratios:

Receiving Assembly Finishing
Square footage
Number of receiving orders

Allocations:

      General Factory Receiving Assembly Finishing
Direct overhead cost $ $ $ $
Allocate:
  General Factory
  Receiving
Total $ $ $ $

2. Using direct labor hours, compute departmental overhead rates. (Round to the nearest cent.)

Overhead Rate
Assembly per direct labor hour
Finishing per direct labor hour
Answer #1

1) Allocation ratios:

Receiving Assembly Finishing
Square footage (2700/13500)= 0.20 (5400/13500)= 0.40 (5400/13500)= 0.40
Number of receiving orders 0 (1680/2700)= 0.62 (1020/2700)= 0.38

Total number of square footage= 2700+5400+5400= 13500

Total number of receiving orders= 1680+1020= 2700

Allocations:

      General Factory Receiving Assembly Finishing
Direct overhead cost $400000 $160000 $43000 $74000
Allocate:
  General Factory (400000) (400000*0.20)= 80000 (400000*0.40)= 160000 (400000*0.40)= 160000
  Receiving - (240000) (240000*0.62)= 148800 (240000*0.38)= 91200
Total $0 $0 $351800 $325200

2. Using direct labor hours, compute departmental overhead rates. (Round to the nearest cent.)

Overhead Rate
Assembly (351800/25000)= $14.07 per direct labor hour
Finishing (325200/40000)= $8.13 per direct labor hour